# Exploratory Causal Inference in SAEnce

### v1

@@ Abstract
Figure 1: Pipeline for Exploratory Causal Inference: (i) Collect data from a Randomized Controlled Trial, (ii) Extract representations via a Foundation Model and Sparse Autoencoder, (iii) Apply Neural Effect Search, and (iv) domain experts interpret the causal findings.
@@ 1 Introduction
In science, data has been historically collected to answer specific questions (Popper, 2005). In this rational view, scientists formulate a hypothesis, often as a causal association, and collect data to falsify it. For example, an experimental ecologist may suspect that exposure to some substance may affect how ants behave, or more in general, “a treatment $T$ has a causal effect on an outcome YY”. They then perform a controlled experiment, administering $T$ or a placebo to a number of individuals and check whether there is a significant difference in the correlation between the treatment assignment and the outcome. While this paradigm has dominated science for centuries, modern science started embracing the creation of atlases: vast, comprehensive maps of natural phenomena, collected for general purposes. Today, we have planetary-scale maps of life genomes (Chikhi et al., 2024), sequencing of 33 different types of cancer (Weinstein et al., 2013), imaging of cells under thousands of perturbations (Sypetkowski et al., 2023) to name a few. Different than the classical paradigm, these datasets call for an empiricist view, starting with exploratory data-driven investigations. The new challenge is that the immense size of these datasets prohibits scientists from just “looking at the data and finding out what is interesting”. Even beyond atlases, consider the specific example of experimental ecology, where fine-grained social interactions between many individuals are critical to understanding the spread of disease (Finn et al., 2019). Clearly, this can be dramatically accelerated with computer vision, using the predictions of a model as input for causal inference pipelines (Cadei et al., 2025). Still, scientists need to decide what to annotate a priori before they can meaningfully look at and understand the data. This introduces a biasing effect, known as the “Matthew effect” (Merton, 1968) or informally as “rich-get-richer”: scientists are biased by prior successful investigations, e.g., behaviors that they have already studied.
@@
In this paper, we characterize differences and synergies between the classical rationalist view and the emerging empiricist one and propose a method to identify statistically significant effects in exploratory experiments, formally grounding it with the language of statistical causality, see Figure 1. We formulate this problem as analyzing a randomized controlled trial, where a treatment is administered randomly and the possible effects are measured indirectly, e.g., via imaging or other raw observations. Instead of scientists formulating only a priori hypotheses on the effect, label some data, and train a model to extend labels to the whole dataset (i.e., the rationalist view (Cadei et al., 2024, 2025)), we propose to train sparse autoencoders (SAEs) on the representation of foundation models and generate data-driven effect hypotheses (i.e., empiricist approach) by the interpretation of the significant effects on the neural representations. In this new paradigm, the main challenge is that, if the SAE is not perfectly disentangled (Elhage et al., 2022), any neuron minimally entangled with the true effect may appear significantly treatment-responsive, which complicates interpretation. To address it, we propose a novel recursive stratification technique to iteratively correct the effect on entangled neurons.
@@
Looking at the data before committing to any hypothesis, we overcome the Matthew effect, enriching the rationalist view in a data-driven way. We propose to work with pretrained foundation models, training SAEs directly on the target experimental data. This is important because pretrained foundation models can be biased as well, which is problematic for drawing scientific conclusions (Cadei et al., 2024). Instead of directly testing a single hypothesis, our approach enables to preliminary explore thousands of potential effects in a semantically expressive latent space, still allowing the domain experts to interpret, judge and eventually test them a posteriori. This is in stark contrast with preliminary empiricist approaches in causality like “causal feature learning” (Chalupka et al., 2017), which only commits to a single hypothesis by discrete clustering. Our contributions are:
@@
  • We propose a novel empiricist methodology building on foundation models and sparse autoencoders. We characterize the statistical challenges in multiple hypothesis testing to discover treatment effects over neural representations in our paradox of exploratory causal inference. Then, we introduce a novel iterative hypothesis testing procedure to overcome such challenges.
  • We showcase in both semi-synthetic (real images but synthetic causal relations) and a real-world trial in experimental ecology that our approach is capable of disentangling and identify the treatment effect in an experiment. To the best of our knowledge, this is the first successful application of sparse autoencoders to causal inference, which we also test in a real-world scientific dataset.
@@ 2 Treatment Effect Estimation in Randomized Controlled Trials
Causal Inference. Causal Inference aims to quantify the effects of an intervention on a certain variable treatment on some outcome variables of interest, see Figure 2 (left). For simplicity, we consider a binary treatment $T=\left\{0,1\right\}$ (e.g., taking a placebo or a drug) and outcome variables $Y\in\left\{0,1\right\}^{r}$ (e.g., binary indicators for symptoms, biomarkers, or clinical events). While continuous extensions would be interesting, we focus on discrete outcomes since continuous concepts in SAEs are not well understood yet (Quirke et al., 2025). At population level we aim to estimate the Average Treatment Effect (ATE):
$\tau=\mathbb{E}[Y(T=1)-Y(T=0)],$ (1)
where $Y(T=1)$ and $Y(T=0)$, or $Y(1)$ and $Y(1)$ for short, denote the potential outcomes under treatment and control (Rubin, 1974) (equivalently $Y|do(T)=1$ and $Y|do(T)=0$ according to Pearl (2009)). Estimating $\tau$ is challenging because, for each individual, only one potential outcome is factually observed—the one under the received treatment—so the counterfactual is missing (fundamental problem of causal inference (Holland, 1986)). This problem is mitigated in the sciences by performing, whenever possible, a Randomized Controlled Trial (RCT). By randomly assigning the treatment, i.e., $T$ has no causes, we prevent spurious correlations between the treatment and any other cause $W\in\mathbb{R}^{q}$ of the outcome (no confounders), allowing to (statistically) identify the ATE with the associational difference, i.e.,
$\tau=\mathbb{E}[Y\mid T=1]-\mathbb{E}[Y\mid T=0],$ (2)
under standard causal assumptions (Rubin, 1986) of consistency (observing $T=t$, then $Y=Y(t)$), and no interference across individuals (i.e., all individuals are independent samples from the population, and the treatment assignment to the individual $i$ does not affect individual $j$). It follows that the difference between the treated and control groups’ sample means is already an unbiased estimator of the ATE. Nonetheless, more sophisticated estimators such as Augmented Inverse Propensity Weighting (AIPW (Robins et al., 1994)) can achieve lower variance and thus greater efficiency.
@@
Figure 2: Exemplary graphical models for randomized controlled trials (i.e., no edge from $W$ to $T$). In Causal Inference (left), both $T$ and $Y$ are observed, and $W$ does not influence $T$ as we are assuming a randomized controlled trial. In Prediction-Powered Causal Inference (center), $Y$ is not observed directly but is known and can be partially labeled. The missing $Y$ is predicted by a neural network from high-dimensional $X$ that is trained either on the same trials if labels are available (Cadei et al., 2024) or on other trials with the same label space (Cadei et al., 2025). In Exploratory Causal Inference (right), $Y$ is unknown and unobserved and is discovered by a neural network from high-dimensional $X$ in a purely data-driven way.
@@
Prediction-Powered Causal Inference and the rationalist approach. Assume that $Y$ is not observed directly. Instead, we measure it indirectly via an high-dimensional variable $X\in\mathcal{X}\subseteq\mathbb{R}^{p}$, capturing the affected outcome information, i.e., $H(Y|X)=0$, mixed with the other attributes of the individual $W\in\mathbb{R}^{q}$. For example, in the trial by Cadei et al. (2024), ants are treated with an invisible substance, which affects their grooming behaviors. Ecologists do not record the behaviors directly but rather take videos $X$ of the ant interactions, which they then analyze. Prior work by  Cadei et al. (2024, 2025) showed how to train a model on partially labeled data or similar experiments to predict factual outcomes $\hat{Y}$ from $X$ that approximate $Y$ and then use them for causal inference. For simplicity, we assume that $T$ is not directly visible in $X$, a common practice in double-blind randomized trials (e.g., neither the patient nor the doctor analyzing the results knows which treatment was assigned). The set-up is illustrated in Figure 2 (center). To simplify the notation, we ignore that some covariates $W$ may only influence $X$ and not $Y$. If such covariates exist, we group them into $W$ (having a null causal effect on $Y$).
@@
Exploratory Causal Inference and the empiricist approach. The rationalist view requires knowing what the treatment will affect a priori, which is also prone to the Matthew effect (Merton, 1968) in exploratory experiments (hypotheses are often informed by the outcome of prior successful trials). In this paper, we consider the setting where experiments are exploratory, which we informally model as the scientists having no a priori knowledge of what $Y$ may be. This is shown in Figure 2 (right), with $Y$ being unobserved and unknown (only measured through $X$). We remark that this problem is related to causal abstraction (Rubenstein et al., 2017; Chalupka et al., 2017). In principle, one may consider the pixels themselves as influenced by the treatment. We instead consider the ground truth $Y$ to be the coarsest possible abstraction of the effect of $T$. In other words, we have that $T\perp\!\!\!\perp W|Y$ and the mutual information $I(Y,X)$ is as small as possible (Achille and Soatto, 2018; Fumero et al., 2023). With a slight abuse of notation, we do not need to assume that such $Y$ exists, so $r$ can be zero if the treatment has no effect at all. Our goal is to propose candidate effects $Y$ to the scientists in a purely data-driven way, discovering significant statistics that differentiate the treated and control populations. It is important to remark that we do not interpret these statistics as necessarily scientifically relevant. The reason is that, when working with high-dimensional data, there can be irrelevant effects, i.e., visible treatment and (finite sample) experiment design biases. Our approach is to identify all significant statistics and leave the interpretation to the domain experts. The empiricist view should not replace the rationalist one, but enrich it with additional data-driven hypotheses.
@@ 3 Exploratory Causal Inference via Neural Representations
To detect treatment effects when only high–dimensional indirect outcome measurements $X$ are available, we turn these raw observations into analyzable measurements. We first pass samples $x$ through a pretrained foundation model (FM) (Bommasani et al., 2022), obtaining representations $h=\phi(x)\in\mathbb{R}^{d}$ whose geometry captures semantically meaningful regularities (Amir et al., 2022; Valeriani et al., 2023). Throughout, we assume the FM is sufficient for the outcome information (Achille and Soatto, 2018) (i.e., $I(X,Y)=I\big(\phi(X),Y\big),$) so working in $h$ preserves exactly the information about the (unknown) outcome factors $Y$ that is present in the raw data. Under sufficiency, any arm difference that exists in $X$ is detectable in representation space, making $h$ a principled surrogate for measurement.
From FM features to a measurement dictionary. While FM features are semantically structured, individual coordinates in $h$ generally not align with human–readable concepts (Bricken et al., 2023). We therefore reparameterize the representation into a sparse, interpretable measurement dictionary using a sparse autoencoder (SAE) (Bricken et al., 2023; Huben et al., 2024). Intuitively, the SAE expresses each $h$ as a sparse linear combination of atoms that behave like measurable channels; sparsity biases solutions toward localized, approximately monosemantic features that scientists can inspect a posteriori. Given foundation model’s features $h\in\mathbb{R}^{d}$, the SAE computes a high–dimensional but sparse code $z\in\mathbb{R}^{d}$ and reconstructs $h$ linearly:
@@
where $\mathbf{E},\mathbf{D}\in\mathbb{R}^{d\times m}$ are respectively the encoder, and decoder linear maps, $b_{e},b_{d}\in\mathbb{R}^{m}$ are the learnable biases, and $g:\mathbb{R}^{m}\!\to\!\mathbb{R}^{m}$ is the encoder nonlinearity (Bricken et al., 2023). Training minimizes a reconstruction loss with a sparsity penalty $\mathcal{S}$ weighted $\lambda\geq 0$, i.e.,
@@
Thereafter, each input can be summarized as ${h}\approx b_{d}+\sum_{j}z_{j}d_{j}$, where $\|z\|_{0}\ll d$ and $\mathbf{D}=[d_{1},\ldots,d_{m}]$. This turns the FM representation into a large dictionary of interpretable channels: each coordinate $z_{j}$ serves as a putative detector of a simple attribute, with still some inevitable leakage (Huben et al., 2024).
Monosemanticity, leakage, and entanglement. In exploratory experiments, we would like each SAE code coordinate to behave like a single, human–readable measurement channel for a simple outcome factor. When this happens, a scientist can read off “what changed” from the few activated codes. In practice, however, codes often leak across factors: one neuron can respond weakly to several distinct attributes, i.e., weak polysemanticity and corresponding entanglement (Locatello et al., 2019). We need a minimal language to talk about (i) the direction in code space associated with a factor and (ii) how widely those directions spill across neurons. Let $Z\in\mathbb{R}^{m}$ be SAE codes and $Y=(Y_{1},\dots,Y_{r})$ the (unknown) binary outcome factors with $m\gg r$. To define the leakage set and index, we first define the concept $Y_{k}$ neural representation as:
@@
Codes as statistical measurement channels. Under FM sufficiency and an (approximately) monosemantic SAE, it becomes natural to pose causal questions at the level of individual codes. If the true affected outcomes $Y$ are perfectly localized in disjoint subsets of coordinates of $Z$, then one can test each coordinate for a treatment–control mean shift using a two–sample tt–test, applying Bonferroni adjustment (Bonferroni, 1936) to control the family–wise error rate at $\alpha$ regardless of the number of tests $m$. This provides an idealized measurement interface: we can scan $Z$ for treatment–responsive channels and later interpret significant coordinates via the dictionary atoms $d_{j}$.

### v2

@@ Abstract
Figure 1: Pipeline for Exploratory Causal Inference: (i) collect experimental data, (ii) extract interpretable representations via a Foundation Model and Sparse Autoencoder, (iii) identify treatment effects via Neural Effect Search, and (iv) interpret the causal findings, i.e., affected neurons, with domain experts.
@@ 1 Introduction
In science, data has been historically collected to answer specific questions (popper2005logic). In this rationalist view, scientists formulate a hypothesis, often as a causal query, and collect data to falsify it. For example, experimental ecologists studying social immunity in ants (cremer2007social) may suspect that exposure to a specific micro-particle may affect specific social behaviors, or more formally, “a treatment $T$ has a causal effect on an outcome behavior YY”. They then perform a controlled experiment, administering the treatment or a placebo to a sample of individuals and check whether there is a significant difference in the correlation between the treatment assignment and the outcome. While this paradigm has dominated science for centuries, modern science has started embracing the creation of atlases: vast, comprehensive maps of natural phenomena, collected for general purposes. Today, we have planetary-scale maps of life genomes (chikhi2024logan), sequencing of 33 different types of cancer (weinstein2013cancer), and imaging of cells under thousands of perturbations (sypetkowski2023rxrx1) to name a few. Different than the classical paradigm, these datasets call for an empiricist view, starting with exploratory data-driven investigations. The new challenge is that the immense size of these datasets prohibits scientists from simply “looking at the data and finding out what is interesting”. Even beyond atlases, in our motivating example from experimental ecology, fine-grained social interactions between many individuals are complex yet critical for understanding the spread of a disease (finn2019use). Hopefully, data processing can be dramatically accelerated with modern machine learning, e.g., relying on pre-trained foundational models and leveraging artificial predictions within causal inference pipelines (cadei2025causal). Still, scientists need to decide what to “look at” a-priori before to validly annotate the data and infer causal relationships, potentially biasing the analyses towards recurrently studied behaviors, phenomena known as the “Matthew effect” (merton1968matthew) or informally as “rich-get-richer”: scientists are biased by prior successful investigations.
Motivating Example: behavioral effects in Social Immunity research [Uncaptioned image] Experimental ecologists collect video data of ants’ interactions under different treatments to study social immunity policies manifested in specific behaviors (cremer2007social; cadei2024smoke). In practice, these behaviors are not directly measured, but annotated from the videos by hand or using machine learning. Which behaviors to study is determined by the scientists’ prior (not data driven), which may not be a priori clear or exhaustive in exploratory experiments.
@@
In this paper, we characterize differences and synergies between the classical rationalist view and the emerging empiricist one and propose a method to identify statistically significant effects in exploratory experiments, formally grounding it with the language of statistical causality, see Figure 1. We formulate the problem on experimental data, where a treatment is administered randomly and the possible effects are measured indirectly, e.g., via imaging or other raw observations. Instead of requiring scientists to formulate treatment effect hypotheses, label some data, and train a model to extend labels to the whole dataset (i.e., the rationalist view (cadei2024smoke; cadei2025causal)), we propose to train a sparse autoencoder (SAE) on the representation of a foundation model and generate data-driven effect hypotheses interpreting the significant effects on the neural representations (i.e., empiricist approach). In this new paradigm, the main challenge is that, if the SAE is not perfectly disentangled (superposition), any neuron minimally entangled with the true effect may appear significantly treatment-responsive, challenging the results interpretation. To address it, we propose a novel recursive stratification technique to iteratively correct the effect on entangled neurons.
@@
Looking at the data before committing to any hypothesis, we overcome the Matthew effect, enriching the rationalist view in a data-driven way. We propose to work with pretrained foundation models, training SAEs directly on the target experimental data. This is important because pretrained foundation models can be biased as well, which is problematic for drawing scientific conclusions (cadei2024smoke). Instead of directly testing a single hypothesis, our approach enables to preliminary explore thousands of potential effects in a semantically expressive latent space, still allowing the domain experts to interpret, judge and eventually test them a posteriori. This is in stark contrast with preliminary empiricist approaches in causality like “causal feature learning” (chalupka2017causal), which only commits to a single hypothesis by discrete clustering. Our contributions are:
@@
  • We propose a novel empiricist methodology building on foundation models and sparse autoencoders. We characterize the statistical challenges in multiple hypothesis testing to discover treatment effects over neural representations in the paradox of exploratory causal inference (ours). Then, we introduce Neural Effect Search, a novel iterative hypothesis testing procedure to overcome such challenges.
  • We showcase in both semi-synthetic (real images with controlled causal relations) and a real-world trial in experimental ecology that our approach is capable of identifying the treatment effects in a exploratory experiment. To the best of our knowledge, this is the first successful application of sparse autoencoders to causal inference, tailored to the analysis of scientific experiments.
@@ 2 Treatment Effect Estimation in Randomized Controlled Trials
Figure 2: Exemplary graphical models for randomized controlled trials to infer the treatment effect (in light gray the observed variables). In classic Causal Inference (left), both the treatment $T$ and outcome of interest $Y$ are observed and unconfounded. In Prediction-Powered Causal Inference (center), $Y$ is only measured indirectly via a complex high-dimensional measurements $X$, partially labeled. The missing $Y$ are retrieved from $X$, training a predictive model on the annotated data. In Exploratory Causal Inference (right), the affected outcome $Y$ is unknown, still measured within $X$, from which a model aims to identify it.
Causal Inference. Causal Inference aims to quantify the effects of an intervention on a certain variable treatment on some outcome variables of interest, see Figure 2 (left). For simplicity, we consider a binary treatment $T=\left\{0,1\right\}$ (e.g., taking a placebo or a drug) and outcome variables $Y\in\left\{0,1\right\}^{r}$ (e.g., binary indicators for symptoms, biomarkers, or clinical events). While continuous extensions would be interesting, we focus on discrete outcomes since continuous concepts in SAEs are not well understood yet (binarySAE). At population level we aim to estimate the Average Treatment Effect (ATE):
$\tau=\mathbb{E}[Y(T=1)-Y(T=0)],$ (1)
where $Y(T=1)$ and $Y(T=0)$, or $Y(1)$ and $Y(1)$ for short, denote the potential outcomes under treatment and control (rubin1974estimating) (equivalently $Y|do(T)=1$ and $Y|do(T)=0$ according to pearl2009causality). Estimating $\tau$ is challenging because, for each individual, only one potential outcome is factually observed—the one under the received treatment—so the counterfactual is missing (fundamental problem of causal inference (holland1986statistics)). This problem is mitigated in the sciences by performing, whenever possible, a Randomized Controlled Trial (RCT). By randomly assigning the treatment (i.e., $T$ has no causes), we prevent spurious correlations between the treatment and any other cause $W\in\mathbb{R}^{q}$ of the outcome (no confounders), allowing to (statistically) identify the ATE with the associational difference, i.e.,
$\tau=\mathbb{E}[Y\mid T=1]-\mathbb{E}[Y\mid T=0],$ (2)
under standard causal assumptions (rubin1986comment) of consistency (observing $T=t$, then $Y=Y(t)$), and no interference across individuals (i.e., all individuals are independent samples from the population, and the treatment assignment to the individual $i$ does not affect individual $j$). It follows that the difference between the treated and control groups’ sample means is already an unbiased estimator of the ATE. Nonetheless, more sophisticated estimators such as Augmented Inverse Propensity Weighting (AIPW (robins1994estimation)) can achieve lower variance and thus greater efficiency.
@@
Prediction-Powered Causal Inference and the rationalist approach. Assume that $Y$ is latent and not observed directly. Instead, we observe it indirectly within high-dimensional measurements $X\in\mathcal{X}\subseteq\mathbb{R}^{p}$, capturing all the affected outcome information, i.e., $H(Y|X)=0$, mixed with the other (latent) outcome causes $W\in\mathbb{R}^{q}$. In our motivating example, let $Y$ representing a well-defined behavior of interest and $X$ its video recording, for which we don’t have a perfect behavior classifier yet. Prior work by  cadei2024smoke; cadei2025causal showed how to train a model on partially labeled data or similar experiments to predict factual outcomes $\hat{Y}\approx Y$ from $X$ and then use them for causal inference. For simplicity, we assume that $T$ is not directly observable in $X$, a common practice in double-blind randomized trials (e.g., neither the patient nor the doctor analyzing the results knows which treatment was assigned). The set-up is illustrated in Figure 2 (center). Both Causal Inference and Prediction-Powered Causal Inference represent rationalist approaches in scientific discovery, ignoring the hypothesis generation preliminary, i.e., data-driven $Y$ definition.
@@
Exploratory Causal Inference and the empiricist approach. The rationalist approaches suffer the Matthew effect (merton1968matthew), narrowing the effect hypotheses on the outcome of prior successful trials. In this paper, we consider the setting where experiments are exploratory, which we informally model as the scientists having no or partial a priori knowledge of what the latent effect may be. This is shown in Figure 2 (right), with $Y$ representing the unobserved and unknown latent factors affected by the treatment T and measure in $X$. In principle, one may consider the pixels themselves as influenced by the treatment. We instead consider the ground truth $Y$ to be the coarsest possible abstraction of the treatment effect in $X$  (rubenstein2017causal; chalupka2017causal). In other words, $[Y,W]$ is the minimal and sufficient representation for $X$ (achille2018emergence; fumero2023leveraging) that renders the treatment independent from the observations, i.e., $T\perp\!\!\!\perp X|[W,Y]$, where only $Y$ is causally affected by a controlled intervention on $T$. With a slight abuse of notation, we do not need to assume that such $Y$ exists, so $r$ can be zero if the treatment has no effect at all. Our goal is to propose candidate effects $Y$ to the scientists in a purely data-driven way, discovering significant statistics that differentiate the treated and control populations. It is important to remark that we do not interpret these statistics as necessarily scientifically relevant. The reason is that, when working with high-dimensional data, there can be irrelevant effects, i.e., visible treatment and (finite sample) experiment design biases. Our approach is to identify all significant statistics and leave the interpretation to the domain experts. The empiricist view should not replace the rationalist one, but enrich it with additional data-driven hypotheses.
@@ 3 Exploratory Causal Inference via Neural Representations
To detect treatment effects when only high–dimensional indirect outcome measurements $X$ are available, we turn these raw observations into analyzable measurements. We first pass samples $x$ through a pretrained foundation model (FM) (fm), obtaining representations $h=\phi(x)\in\mathbb{R}^{d}$ whose geometry captures semantically meaningful regularities (visualmeaning; languagemeaning). Throughout, we assume the FM is sufficient for the outcome information (achille2018emergence), i.e., $I(X,Y)=I\big(\phi(X),Y\big)$, so working in $h$ preserves exactly the information about the (unknown) outcome factors $Y$ that is present in the raw data. Under sufficiency, any arm difference that exists in $X$ is detectable in representation space, making $h$ a principled surrogate for measurement.
From FM features to a measurement dictionary. While FM features are semantically structured, individual coordinates in $h$ generally not align with human–readable concepts (bricken2023monosemanticity). We therefore reparameterize the representation into a sparse, interpretable measurement dictionary using a sparse autoencoder (SAE) (bricken2023monosemanticity; huben2024sparse). Intuitively, the SAE expresses each $h$ as a sparse linear combination of atoms that behave like measurable channels; sparsity biases solutions toward localized, approximately monosemantic features that scientists can inspect a posteriori. Given foundation model’s features $h\in\mathbb{R}^{d}$, the SAE computes a high–dimensional but sparse code $z\in\mathbb{R}^{d}$ and reconstructs $h$ linearly:
@@
where $\mathbf{E},\mathbf{D}\in\mathbb{R}^{d\times m}$ are respectively the encoder, and decoder linear maps, $b_{e},b_{d}\in\mathbb{R}^{m}$ are the learnable biases, and $g:\mathbb{R}^{m}\!\to\!\mathbb{R}^{m}$ is the encoder nonlinearity (bricken2023monosemanticity). Training minimizes a reconstruction loss with a sparsity penalty $\mathcal{S}$ weighted $\lambda\geq 0$, i.e.,
@@
Thereafter, each input can be summarized as ${h}\approx b_{d}+\sum_{j}z_{j}d_{j}$, where $\|z\|_{0}\ll d$ and $\mathbf{D}=[d_{1},\ldots,d_{m}]$. This turns the FM representation into a large dictionary of interpretable channels: each coordinate $z_{j}$ serves as a putative detector of a simple attribute, with still some inevitable leakage (huben2024sparse).
Monosemanticity, leakage, and entanglement. In exploratory experiments, we would like each SAE code coordinate to behave like a single, human–readable measurement channel for a simple outcome factor. When this happens, a scientist can read off “what changed” from the few activated codes. In practice, however, codes often leak across factors: one neuron can respond weakly to several distinct attributes, i.e., weak polysemanticity and corresponding entanglement (locatello2019challengingcommonassumptionsunsupervised). We need a minimal language to talk about (i) the direction in code space associated with a factor and (ii) how widely those directions spill across neurons. Let $Z\in\mathbb{R}^{m}$ be SAE codes and $Y=(Y_{1},\dots,Y_{r})$ the (unknown) binary outcome factors with $m\gg r$. To define the leakage set and index, we first define the concept $Y_{k}$ neural representation as:
@@
Codes as statistical measurement channels. Under FM sufficiency and an (approximately) monosemantic SAE, it becomes natural to pose causal questions at the level of individual codes. If the true affected outcomes $Y$ are perfectly localized in disjoint subsets of coordinates of $Z$, then one can test each coordinate for a treatment–control mean shift using a two–sample tt–test, applying Bonferroni adjustment (bonferroni1936teoria) to control the family–wise error rate at $\alpha$ regardless of the number of tests $m$. This provides an idealized measurement interface: we can scan $Z$ for treatment–responsive channels and later interpret significant coordinates via the dictionary atoms $d_{j}$.